### v1

# Counterfactual Tuning for Temporal Sensitivity Enhancement in Large Language Model-based Recommendation

Abstract.
Recent advances have applied large language models (LLMs) to sequential recommendation, leveraging their pre‑training knowledge and reasoning capabilities to provide more personalized user experiences. However, existing LLM‑based methods fail to sufficiently leverage the rich temporal information inherent in users’ historical interaction sequences, stemming from fundamental architectural constraints: LLMs process information through self-attention mechanisms that lack inherent sequence ordering and rely on position embeddings designed primarily for natural language rather than user interaction sequences. This limitation significantly impairs their ability to capture the evolution of user preferences over time and predict future interests accurately.
To address this critical gap, we propose Counterfactual Enhanced Temporal Framework for LLM‑Based Recommendation (CETRec). CETRec is grounded in causal inference principles, which allow it to isolate and measure the specific impact of temporal information on recommendation outcomes. By conceptualizing temporal order as an independent causal factor distinct from item content, we can quantify its unique contribution through counterfactual reasoning—comparing what recommendations would be made with and without temporal information while keeping all other factors constant. This causal framing enables CETRec to design a novel counterfactual tuning objective that directly optimizes the model’s temporal sensitivity, teaching LLMs to recognize both absolute timestamps and relative ordering patterns in user histories. Combined with our counterfactual tuning task derived from causal analysis, CETRec effectively enhances LLMs’ awareness of both absolute order (how recently items were interacted with) and relative order (the sequential relationships between items). Extensive experiments on real-world datasets demonstrate that CETRec consistently outperforms state-of-the-art baselines.
Sequential Recommendation, LLM-based Recommendations, Counterfactual Inference
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Information systems Recommender System
1. Introduction
In the era of information explosion, recommendation systems have become essential tools that help users navigate vast amounts of content across online platforms (Fan et al., 2023). Within this domain, sequential recommendation has gained increasing attention for its ability to effectively model users’ temporal interaction sequences and capture evolving preferences over time, producing more personalized and accurate recommendations (Wu et al., 2023; Tang and Wang, 2018; Hidasi et al., 2016; Kang and McAuley, 2018). Recently, with the rapid advancement of large language models (LLMs), numerous studies have begun to incorporate these powerful models into sequential recommendation systems, leveraging their rich pre-trained knowledge and reasoning capabilities (Bao et al., 2023a; Zheng et al., 2024; Liao et al., 2024; Geng et al., 2022; Bao et al., 2023b; Wei et al., 2024). Clearly, temporal information—the chronological ordering and timing of user interactions—plays a crucial role in sequential recommendation, as it directly reflects the evolution of user interest and preference dynamics.
Figure 1. Illustration of absolute vs. relative order effects. Users A and B have the same history movie set but differ in absolute order; Users C and D have the same history movie set but differ in relative order.
[…]
Different user interest shown in different interaction sequences.
In sequential recommendation, user historical interactions encompass two aspects of temporal information: absolute order information and relative order information. Absolute order refers to the actual timestamps of interactions (i.e., how long ago each event occurred), whereas relative order refers to the ordinal relationships among users’ multiple interactions. Both aspects play a crucial role in predicting the next item a user may interact with. For example, as shown in Figure 1, user A who watched Iron Man and Iron Man 2 within the last month is significantly more likely to proceed to Iron Man 3 than user B whose viewing occurred ten years ago. Moreover, if user D watches several romance films immediately after Iron Man and Iron Man 2, their probability of watching Iron Man 3 is typically lower than that of user C who views those same romance films before watching the first two Iron Man installments, because the former’s interests may already have shifted.
Existing LLM-based recommendation approaches can be largely grouped into two main streams: 1) Paradigm adaptation methods that attempt to reformulate recommendation tasks into formats suitable for LLMs’ text understanding and reasoning capabilities (Bao et al., 2023a; Geng et al., 2022). These approaches typically transform recommendation into natural language generation tasks, framing item suggestions as text completion or response generation problems. They leverage prompting techniques and fine-tuning strategies to align LLMs’ pre-trained knowledge with recommendation objectives. 2) Multimodal integration methods treat items as a distinct data modality (Yang et al., 2023; Lin et al., 2023). These approaches construct specialized item embeddings that capture item characteristics, user preferences, and interaction patterns, then integrate these representations with LLMs’ text understanding capabilities through cross-modal attention mechanisms or embedding alignment techniques, leveraging LLMs’ rich pre-trained knowledge to enhance representation of items. Despite these advancements, these approaches primarily focus on modeling the semantic aspects of interacted items or optimizing the overall recommendation task, and do not explicitly capture the temporal structure of user interaction sequences. This temporal oversight leaves a significant gap in their ability to model users’ evolving interests and time-dependent preference shifts. An empirical study (Kim et al., 2025) that perturbs interaction sequences shows that LLM-based recommenders exhibit smaller performance drops under different sequence orders compared to traditional models, further demonstrating their limited ability to capture temporal information.
Large language models, built on Transformer architectures, fundamentally rely on the self-attention mechanism which inherently lacks any built-in notion of sequence order (Vaswani et al., 2017). While these models incorporate position embeddings as an external means of injecting token order for natural language processing tasks, such mechanisms prove inadequate for capturing the complex temporal dynamics in sequential recommendation. This inadequacy stems from several factors: First, position embeddings in LLMs are designed primarily for natural language, where grammatical structure and semantic context provide additional ordering cues that are absent in item interaction sequences. Second, these embeddings are typically applied at the token level rather than the item level, preventing the model from treating each interacted item as a cohesive temporal unit during inference. Third, recommendation prompts typically include extensive task instructions and contextual background information, causing the already sparse temporal signals to become ”diluted” amid multiple sources of information. As a result, the model struggles to correctly interpret the temporal relationships between user interactions, limiting its ability to capture the dynamic evolution of user preferences that is essential for accurate sequential recommendation. This fundamental limitation suggests that standard position embeddings in LLMs, while sufficient for natural language tasks, cannot adequately capture the nuanced temporal relationships crucial for effective sequential recommendation.
To address this challenge, we propose Counterfactual Enhanced Temporal Framework for LLM‑Based Recommendation (CETRec), which leverages causal inference to explicitly model and enhance temporal sensitivity in recommendation. CETRec is grounded in a formal causal model that distinguishes temporal order as a distinct causal factor from the items themselves in determining recommendation outcomes. This causal framing allows us to isolate and measure the specific impact of temporal information through counterfactual reasoning. To model this temporal information, for each item, we assign every token the same additional position embedding on top of its original positional embedding, called temporal embedding. Subsequently, by constructing a counterfactual world where all temporal order information is erased, specifically, replacing each item’s temporal embedding with that of the first item, while keeping all other factors constant, we can precisely quantify the causal effect of temporal order on recommendations. Based on this counterfactual inference, we design a novel tuning objective that maximizes the divergence between predictions in the factual (temporally ordered) and counterfactual (temporally erased) worlds. This approach directly optimizes the model’s sensitivity to temporal patterns by encouraging it to produce different predictions when temporal information is altered, effectively teaching the model to recognize and utilize both absolute and relative temporal signals in user interaction sequences.
In summary, our contributions are threefold:
  • We identify and analyze the insufficient use of temporal information in user interaction histories by existing LLM-based recommenders in the sequential recommendation setting. Through systematic analysis, we reveal how architectural limitations in Transformer models and position embedding schemes designed for natural language processing create fundamental barriers to capturing temporal dynamics in user preference evolution.
  • We propose a novel framework, Counterfactual Enhanced Tempo-ral Framework for LLM‑Based Recommendation (CETRec), which employs a principled causal approach to recommendation. CETRec integrates temporal embedding at the item level and introduces a counterfactual tuning task derived from causal inference that directly optimizes the model’s sensitivity to both absolute and relative temporal ordering. By formalizing temporal order as an independent causal factor in our structural causal graph, we enable precise measurement and enhancement of its contribution to recommendation outcomes.
  • We conduct extensive experiments across multiple real-world datasets, validating the effectiveness of CETRec in enhancing temporal information comprehension. Through ablation analysis, we further verify CETRec’s superior ability to leverage temporal patterns in generating recommendations that align with users’ evolving interests.
2. Related Work
2.1. LLM-based Recommendation
Sequential recommendation predicts the next item a user may be interested in by leveraging user and item features extracted from the user’s historical interaction sequence. Early methods relied on Markov chains to capture transition probabilities between items (Rendle et al., 2010). With the advent of deep learning, a variety of architectures—RNN (Hidasi et al., 2016; Cui et al., 2020; Beutel et al., 2018), CNN (Tang and Wang, 2018; Yuan et al., 2019; Yan et al., 2019) and attention (Kang and McAuley, 2018; Zhang et al., 2019)—have been employed to model user–item interaction dynamics.
Recently, large language models (LLMs) have achieved remarkable success across domains (Zhao et al., 2023; Chowdhery et al., 2023; Wei et al., 2022; Driess et al., 2023; Ai et al., 2023; Jin et al., 2023) thanks to two key strengths: their rich pre-training knowledge and powerful reasoning capabilities. Following these strengths, LLM‐based approaches to sequential recommendation broadly fall into two paradigms:
  • Representation Enhancement. These methods leverage the LLMs’ pre-trained knowledge to improve item or user representations—either by extracting embeddings from items’ textual (or other modal) representations (Yang et al., 2023; Lin et al., 2024; Tan et al., 2024; Wang et al., 2024), or by distilling embeddings from traditional models into the LLMs (Liao et al., 2024; Zheng et al., 2024; Kim et al., 2025).
  • Paradigm Adaption. These methods leverage the reasoning capabilities of LLMs by aligning them with the recommendation paradigm—reformulated as natural language generation tasks—through techniques such as fine-tuning (Geng et al., 2022; Bao et al., 2023a, b; Zhang et al., 2024a), prompting (Dai et al., 2023), or in-context learning (ICL) (Hou et al., 2024).
Although a prior study (Kim et al., 2025) has noted that LLMs inadequately exploit the temporal ordering in user interaction sequences—attempting to distill sequence information from conventional models—the majority of existing methods still lack explicit temporal modeling within the LLM itself. To our best knowledge, we are the first to harness LLMs’ native capacity to directly model the temporal order of interaction sequence.
2.2. Causal Inference for Recommendation
In recommender systems, non‐causal approaches typically model only correlations between variables, which can lead to spurious relationships and limited generalization. In contrast, causal inference aims to reveal the underlying causal mechanisms in the data, enabling more robust and interpretable recommendations (Luo et al., 2023).
Prior work in causal recommendation has primarily focused on three key areas: First, debiasing methods (Zhu et al., 2020; Christakopoulou et al., 2020) employ causal techniques to mitigate various forms of systematic bias in recommendation data. These include popularity bias (Craswell et al., 2008), where frequently interacted items receive disproportionate exposure regardless of actual user preference, and click bias (Wang et al., 2021), where user interaction patterns are influenced by item position rather than genuine interest. By modeling these biases as confounding factors in a causal graph, researchers have developed intervention strategies that produce more balanced and fair recommendations.
Second, causal approaches have proven effective for handling data sparsity and noise (Xiong et al., 2021; Zhang et al., 2021; Liu et al., 2022). By leveraging structural causal models to distinguish genuine user preferences from random interactions or missing data patterns, these methods enhance recommendation robustness in scenarios with limited observations. The causal perspective helps identify which missing interactions represent true negative preferences versus those that are merely unobserved positive preferences, leading to more accurate user modeling.
Third, causal recommendation has advanced objectives beyond accuracy, including explainability (Wachter et al., 2017; Attolou et al., 2024) and fairness (Huang et al., 2022; Zhu et al., 2023). Explainable recommendation benefits from causal reasoning by identifying the actual factors that influence user decisions rather than merely correlative features. Similarly, fairness-aware recommenders use causal techniques to detect and mitigate discrimination against protected groups by modeling the causal relationships between sensitive attributes and recommendation outcomes.
Unlike these previous methods, our approach introduces a novel application of causal inference to sequential recommendation by specifically targeting temporal dynamics. We leverage counterfactual inference to isolate and quantify the causal effect of temporal order on recommendation outcomes, enabling LLMs to better capture the evolution of user preferences over time. By constructing a counterfactual world where temporal information is erased while preserving item content, we can directly measure temporal sensitivity and design specialized tuning objectives that enhance the model’s awareness of both absolute and relative temporal patterns in user interaction sequences.
3. Preliminaries
In this section, we formally define our task and introduce the basic conceptions used in CETRec, including position embedding and causal model.
Table 1. A tuning template for the Game dataset. ¡His_Seq¿ and ¡Target_Item¿ denote the user’s historical interaction sequence and the ground-truth next item, respectively.
Instruction Input
Task Instruction: Given a list of video games the user has played before, please recommend a new video game that the user likes to the user.
[…]
Instruction Output
Output: <Target_Item>
3.1. Task Formulation
In sequential recommendation scenario, the model predicts the next item a user is likely to interact with based on their historical interaction sequence. Formally, let =$\mathcal{I}=\{i_{1},i_{2},...,i_{|\mathcal{I}|}\}$ denote the set of items. Given a temporally ordered interaction sequence $\mathcal{S}=(i_{t_{1}},i_{t_{2}},...,i_{t_{|\mathcal{S}|}})$, where $i_{t_{n}}\in\mathcal{I}$ denotes the item interacted with at time $t_{n}$, the goal of the model is to predict the next item $i_{t_{n}+1}$.
Figure 2. The framework of CETRec. The user interaction sequence is transformed to the instruction tuning format. CETRec constructs temporally ordered and erased sequence through assigning the corresponding position embedding to every token of each item. During counterfactual tuning, CETRec yield Time-aware Loss from temporal oredered sequence and counterfactual tuning loss from the difference between temporally ordered and erased sequence.
[…]
(1) $L_{o}=\sum_{(x,y)\in D}\sum_{k=1}^{|y|}\ell\left(f_{\Theta}(x,y_{<k});\ y_{k}\right),$
where $\ell$ denotes the loss function, $f_{\Theta}(\cdot)$ denotes the LLM with parameters $\Theta$, $y_{k}$ refers to the $k$-th token in $y$, and $y_{<k}$ represents tokens before $y_{k}$.
However, fine-tuning LLMs with full parameters is both time-consuming and computationally expensive. To address this, we adopt LoRA (Low-Rank Adaptation) (Hu et al., 2022), which freezes the original model weights and trains only a small set of additional low-rank matrices, achieving comparable performance with a fraction of the overhead. Formally, The new loss is formulated as:
(2) $L_{n}=\sum_{(x,y)\in D}\sum_{k=1}^{|y|}\ell\left(f_{\Theta+\Phi}(x,y_{<k});\ y_{k}\right),$
where $\Phi$ is the LoRA parameters.
3.2. Position Embedding
Position Embedding(PE) is a method used in Transformer-based models to encode the order of input tokens. Existing position embedding can be broadly categorized into two types: absolute and relative (Dufter et al., 2022). In absolute position embedding, each input token embedding $x_{k}$ at position $k$ is augmented by a learnable or fixed vector $p_{k}$ that depends solely on $k$, typically of the form $x_{k}+p_{k}$. In relative position embedding, position information is incorporated dynamically within the attention mechanism by considering the relative distances between tokens, instead of adding fixed positional signals to the input embeddings.
Sinusoidal Position Embedding(SinPE) is a classic form of absolute position embedding proposed in the original Transformer architecture (Vaswani et al., 2017), defined as:
(3) $\begin{cases}p_{k,2t}&=\sin\left(\frac{k}{10000^{2t/d}}\right)\\ p_{k,2t+1}&=\cos\left(\frac{k}{10000^{2t/d}}\right)\end{cases},$
where $p_{k,2t}$ is the $2t$-th element of the d-dimensional vector $p_{k}$.
Rotary Position Embedding(RoPE) (Su et al., 2024) is a state-of-the-art relative position embedding method that incorporates characteristics of absolute position embedding. It injects positional information by rotating the query ($q$) and key ($k$) vectors during the attention computation based on their positions. The formulation is as follows:
(4) $f_{\{q,k\}}(x_{m},m)=R_{\Theta,m}^{d}W_{\{q,k\}}x_{m},$
where $R_{\Theta,m}^{d}$ is A block-diagonal $d\times d$ matrix composed of $d/2$ independent $2\times 2$ rotation blocks, with $\Theta$ denoting rotary parameters associated with the position $m$ and each corresponding embedding dimension. Applying RoPE to self-attention process, we have:
(5) $q_{m}^{\top}k_{n}=(R_{\Theta,m}^{d}W_{q}x_{m})^{\top}(R_{\Theta,n}^{d}W_{k}x_{n})=x_{m}^{\top}W_{q}R_{\Theta,n-m}^{d}W_{k}x_{n},$
where $R_{\Theta,n-m}^{d}=(R_{\Theta,m}^{d})^{\top}R_{\Theta,n}^{d}$ ,which only depends on the relative distances between tokens. In this work, we employ the two position embedding schemes described above: SinPE and RoPE.
3.3. Causal model.
A causal model (Pearl, 2009) formally describes the relationships between variables through a structured framework, typically defined as a triple: M = ⟨U, V, F ⟩, where:
  • $U$ is the set of exogenous variables determined by factors outside the model;
  • $V$ is the set of endogenous variables determined within the model;
  • $F$ is the set of structural functions such that each $f_{i}\in F$ maps $U_{i}\cup\mathrm{PA}_{i}$ to $V_{i}$, where $U_{i}\subseteq U$, $\mathrm{PA}_{i}\subseteq V\setminus V_{i}$, and the complete set $F$ defines a mapping from $U$ to $V$.
Every causal model $M$ is associated with a causal graph $G(M)$, a directed graph where nodes represent variables, and edges are drawn from $U_{i}$ and $\mathrm{PA}_{i}$ to $V_{i}$.
Figure 3. (a) Causal graph for sequential recommendation task. (b) The twin network of counterfactual inference on temporal order $T$.
[…]
Causal Anal
4. Methodology
In this section, we present a comprehensive framework for enhancing temporal sensitivity in LLM-based sequential recommendation. We first conduct a causal analysis of sequential recommendation from a temporal perspective, establishing the theoretical foundation for our approach. Next, we introduce our novel method for modeling temporal order with position embeddings. We then apply counterfactual inference to quantify the causal effect of temporal information, which leads to the design of our CETRec framework, as illustrated in Figure 2.
4.1. Causal Analysis
Sequential recommendation inherently involves complex temporal dynamics that influence user preferences. To systematically analyze these dynamics, we treat temporal order as an independent variable that, together with the interacted items themselves, defines the user’s historical interaction sequence. From this perspective, we perform a formal causal analysis of the sequential recommendation task and derive the causal graph shown in Figure 3(a). This graph captures the key variables and their interdependencies in the recommendation process. The semantics of the nodes are as follows:
  • Node $I$: items in the user’s interaction history, representing the content aspects of historical interactions regardless of their temporal arrangement.
  • Node $T$: temporal order of interacted items, encompassing both absolute timestamps and relative sequential relationships.
  • Node $H$: the user’s entire historical interaction sequence, which is jointly determined by both the items and their temporal ordering.
  • Node $U$: background or contextual information affecting the recommendation results, such as shifts in user preference, changes in item popularity, or variations in item price/reputation over time.
  • Node $Y$: he next item predicted or generated by the recommender system, which is the target variable we aim to optimize.
This causal graph provides several important insights. Previous studies have primarily focused on the path $H\rightarrow Y$ or path $I\rightarrow H\rightarrow Y$, corresponding to paradigm adaptation methods and multi-modal integration methods, respectively. The former treats recommendation as a sequence-to-item generation task, while the latter enriches item representations with specialized embeddings. However, the influence of temporal order—captured by the path $T\rightarrow H\rightarrow Y$-has been largely overlooked despite its crucial role in shaping user preferences and interaction patterns.
By explicitly modeling temporal order as a distinct causal factor, we can isolate its specific contribution to recommendation outcomes, allowing us to design targeted interventions that enhance LLMs’ temporal sensitivity. This causal framing represents a significant departure from previous approaches that implicitly blend temporal information with content information or treat it as a secondary aspect of the recommendation process.
4.2. Modeling Temporal Order
To effectively encode temporal information in user interaction sequences, we introduce temporal embedding—a cross-token level position embedding approach that preserves the semantic integrity of each item while explicitly representing its position in the user’s interaction history. Specifically, for sequence $\mathcal{S}$, we assign each item $i_{k}$ a position embedding $p_{k}$ based on its position $k$ in interaction sequence, and add $p_{k}$ to every token of that item. Crucially, rather than applying this embedding at the token level (as in traditional Transformer models), we add the same position embedding $p_{k}$ to every token that constitutes the item. This approach treats each item as a cohesive temporal unit, reinforcing its semantic boundaries within the sequence. We also retain the original token-level positional embeddings to preserve the model’s ability to understand contextual relationships within the prompt. Consequently, the final input embedding for the t-th token is given by:
(6) $I_{t}=x_{t}+p_{t}+p_{k},$
where $x_{t}$ and $p_{t}$ denote the token embedding and position embedding of the $t$-th token respectively, and $p_{k}$ represents the temporal embedding of the $k$-th item. Besides, $p_{k}$ is set to null if the $t$-th token does not belong to any item. For example, suppose that the movie Shawshank Redemption appears in the third position of $\mathcal{S}$, tokenized into 5 tokens (i.e., ‘Sh’, ‘aw’, ‘sh’, ‘ank’, ‘Redemption’), then each of these 5 tokens will be augmented by the embedding $p_{3}$, signifying that they collectively represent an item that appeared third in the user’s interaction history. This consistent application of temporal embeddings across all tokens of an item ensures that the model perceives the item as a unified entity within the temporal sequence.
This temporal embedding approach offers two primary advantages over standard position embeddings: 1) Item-level semantic coherence: By applying the same temporal embedding to all tokens of an item, we enable the model to treat each item as a cohesive unit during representation learning and inference. This alignment between semantic boundaries and temporal encoding helps the model better leverage item-level prior knowledge and understand items as integral units in the user’s preference evolution. 2)Explicit temporal encoding: By assigning incrementally increasing position embeddings ($p_{1},p_{2},p_{3},\ldots$) to items based on their interaction order, we explicitly encode both absolute and relative temporal relationships. This makes the temporal structure of the interaction sequence more salient to the model, strengthening its ability to detect patterns like recent interest shifts or consistent preferences over time.
4.3. Counterfactual Inference
Having established our method for encoding temporal information, we now introduce a counterfactual inference framework to enhance the model’s sensitivity to this information. Counterfactual inference allows us to answer questions about what would happen under hypothetical conditions that differ from observed reality.
In our context, we perform counterfactual inference by altering the temporal order variable $T$ to construct a counterfactual world where all temporal information is erased, while preserving all other aspects of the recommendation scenario. This approach enables us to isolate and measure the specific impact of temporal ordering on recommendation outcomes. The counterfactual inference framework evaluates how descendant variables would change under a different treatment variable while holding other conditions constant(Fenton et al., 2020; Pearl, 2001). In our causal graph, this can be expressed as: ‘what $Y$ would be in situation $U=u$ and $I=i$, had $T$ been $t^{*}$?’. This counterfactual query examines an alternative course of history where the temporal structure of interactions is altered, while keeping all other factors fixed.
Traditional methods for estimating counterfactual effects require computing and storing posterior information for all background variables in the model—a requirement that is often impractical for complex systems. To address this challenge, we employ the twin network approach, a graphical method that uses two structurally identical networks to jointly represent the factual and counterfactual worlds (Balke and Pearl, 1994). When performing counterfactual inference on a variable, all arrows entering this variable in the counterfactual world are removed to reflect the intervention, allowing the counterfactual query to be computed using standard probabilistic inference techniques.
In our model, the resulting twin network of our counterfactual inference is shown in Figure 3(b). The counterfactual intervention effectively places all user interactions at the same—and earliest—time step, maximally removing both relative and absolute order information. Concretely, we implement this counterfactual world by replacing each item’s temporal embedding in the input sequence with the temporal embedding of the first item $i_{0}$, denoted $p_{0}$. This intervention creates a scenario where all items appear to have occurred at the same point in time, effectively erasing any sequential patterns or temporal dynamics that might influence recommendation outcomes. The resulting counterfactual world thus preserves the exact set of interacted items and all other latent factors, differing only in its lack of temporal structure.
Based on the twin network framework, the causal effect of temporal order can be computed as the difference between the factual and counterfactual predictions:
(7) $CE_{T}=P(Y\ |\ T=t,I=i,U=u)-P(Y^{*}\ |\ T=t^{*},I=i,U=u),$
where $t$ denotes the original temporal order, while $t^{*}$ represents the erased (counterfactual) order. This causal effect quantifies how much the recommendation outcomes change when temporal information is removed, providing a direct measure of the model’s sensitivity to temporal patterns.
4.4. Model Paradigm
Building on the causal analysis and counterfactual inference framework described above, we now present the complete CETRec model paradigm for both the tuning phase and inference phase.
Tuning. The core insight of CETRec is that we can enhance a model’s temporal sensitivity by explicitly optimizing for the divergence between predictions made with and without temporal information. Based on the causal effect estimated via counterfactual inference, the influence of temporal order on the model’s token generation probability can be expressed as: f_Θ(x_t, y_¡k)  -  f_Θ(x_0, y_¡k), where $f_{\Theta}$ represents the model’s probability distribution over the next token, $x_{t}$ denotes the instruction input with temporally ordered sequence, $x_{0}$ denotes the same input with the temporally erased sequence (the counterfactual world), and $y_{<k}$ represents the generated tokens up to position $k$.
To encourage the model to become more sensitive to temporal information, we design a novel counterfactual tuning loss $L_{CT}$ that directly optimizes this influence:
(8) $L_{CT}=\sum_{(x,y)\in\mathcal{D}}\sum_{k=1}^{|y|}\ell\left(f_{\Theta+\Phi}(x_{t},y_{<k})-f_{\Theta+\Phi}(x_{0},y_{<k});y_{k}\right).$
This loss function encourages the model to produce different predictions when temporal information is altered, effectively teaching it to recognize and utilize both absolute and relative temporal signals in user interaction sequences. By maximizing the divergence between predictions in the factual and counterfactual worlds, we directly optimize the model’s temporal sensitivity.
In addition to the counterfactual tuning loss, we also define a temporal-aware loss $L_{TA}$that enhances the model’s overall recommendation capability using the temporally ordered input:
(9) $L_{TA}=\sum_{(x,y)\in D}\sum_{k=1}^{|y|}\ell\left(f_{\Theta+\Phi}(x,y_{<k});\ y_{k}\right).$
This loss function ensures that the model maintains high recommendation accuracy while developing increased sensitivity to temporal patterns. By combining these two objectives, we create a balanced training regime that improves temporal awareness without sacrificing overall recommendation quality.
The formal objective of CETRec is thus defined as a weighted combination of these two loss functions:
(10) $L=L_{TA}+\lambda L_{CT}$
where $\lambda$ controls the importance of counterfactual loss. This weighting allows us to adjust the trade-off between optimizing for recommendation accuracy and temporal sensitivity, enabling flexible adaptation to different recommendation scenarios and user preference patterns.
Inference. During the inference phase, the model generates predicted items in natural language format based on the instruction input, using the temporally ordered sequence with original temporal embeddings. This approach ensures that the model leverages its enhanced temporal sensitivity to produce recommendations that account for both the content of previous interactions and their temporal arrangement. By preserving the temporal structure during inference, CETRec can effectively capture evolving user preferences, recent interest shifts, and time-dependent patterns in interaction sequences. This temporal awareness enables more personalized and contextually relevant recommendations that align with users’ current interests rather than outdated preferences.
4.5. Discussion
Our methodological design for CETRec highlights several key considerations in temporal modeling for LLM-based recommendation. The temporal embedding approach strikes a balance between preserving language understanding capabilities and enhancing temporal awareness by treating items as cohesive units. The observed differences between SinPE and RoPE implementations—where RoPE achieves better recommendation performance while SinPE shows higher temporal sensitivity—reflect their distinct mathematical properties: RoPE excels at capturing nuanced relative relationships, while SinPE creates more pronounced absolute position signals. Furthermore, our counterfactual inference framework connects to broader trends in causal machine learning, offering advantages over purely correlational methods by explicitly isolating temporal effects.
Table 2. Statistical details of training datasets.
Dataset MovieLens Steam
#Sequences 192,483 151,056
#Items 3,952 3,581
#Interactions 999,611 239,796
5. Experiment
In this section, we conduct experiments to answer the following research questions:
  • RQ1: How does CETRec perform compared to other traditional and LLM-based recommenders?
  • RQ2: How does different position embedding emthods applied in temporal embedding perform?
  • RQ3: What are the respective contributions of the two components of CETRec to its performance?
  • RQ4: How effective is CETRec at enhancing an LLM’s sensitivity to temporal information?
Table 3. Performance of baselines and CETRec on MovieLens and Steam datasets.
MovieLens Steam
HR@5 NDCG@5 HR@10 NDCG@10 HR@5 NDCG@5 HR@10 NDCG@10
Caser 0.0151 0.0126 0.0310 0.0172 0.0355 0.0315 0.0529 0.0367
GRU4Rec 0.0335 0.0279 0.0511 0.0329 0.0397 0.0362 0.0538 0.0402
SASRec 0.0470 0.0265 0.0797 0.0371 0.0544 0.0376 0.0653 0.0411
P5 0.0538 0.0334 0.0911 0.0452 0.0688 0.0468 0.1049 0.0583
BIGRec 0.0985 0.0614 0.1093 0.0650 0.0595 0.0408 0.0687 0.0438
CFT 0.1200 0.0733 0.1299 0.0765 0.0737 0.0498 0.0838 0.0530
CETRec (SinPE) 0.1101 0.0693 0.1209 0.0728 0.0670 0.0436 0.0779 0.0471
CETRec (RoPE) 0.1283 0.0779 0.1365 0.0806 0.0888 0.0635 0.0997 0.0670
5.1. Experiment Settings
5.1.1. Datasets
We conduct experiments on two real real-wrold datasets: MovieLens [^1] https://grouplens.org/datasets/movielens/ and Steam (Kang and McAuley, 2018). MovieLens dataset is a well-known movie recommendation dataset containing the movies titles, user ratings and timestamps. Steam dataset is a video-games recommendation dataset that records the title of games users have played on the Steam Store. In both datasets, user interactions can be organized into sequences in chronological order.
Due to the massive number of parameters in large language models, tuning an LLM consumes substantial time and computational resources, so the dataset size must be kept within a reasonable range. Therefore, for MovieLens dataset, we choose MovieLens1M dataset and sample $1/10$ of interactions. For the Steam dataset, we first retained users who had played at least 20 games, and then randomly sampled $1/3$ of users and $1/3$ of games. For the two datasets, we construct, interaction sequences of length 3 to 10 based on each user’s item interactions. Then we sort all sequences in chronological order and then split the data into training, validation, and testing data at the ratio of 8:1:1. This strategy ensures that future interactions are excluded from the training set, thereby preventing information leakage. The statistical details of training datasets are summerized in Table 2.
5.1.2. Evaluation Settings
For each sequence, we designate the item immediately interacted after it as the target. We employ all-ranking protocol for evaluation, treating all non-interacted items as potential candidates and select two widely used metrics: Hit Ratio (HR@K) and Normalized Discounted Cumulative Gain (NDCG@K) (Kang and McAuley, 2018; Bao et al., 2023a), with $K\in\{5,10\}$.
5.1.3. Baselines
We demonstrate the ability of CETRec by comparing it to the following traditional and LLM-based methods:
  • GRU4Rec (Hidasi et al., 2016). An RNN-based recommender that uses Gated Recurrent Units (GRUs) to encode user interaction histories. It effectively captures both short- and long-term dependencies in sequential click data.
  • Caser (Tang and Wang, 2018). A CNN-based model that represents the last $L$ interactions as an $L\times d$ embedding “image” and applies horizontal and vertical convolutional filters. It learns both local item transition patterns and global user preferences for next-item prediction.
  • SASRec (Kang and McAuley, 2018). A Transformer-based recommender that employs masked multi-head self-attention to attend to all previous items in a sequence. It excels at modeling complex dependencies without relying on recurrent connections.
  • P5 (Geng et al., 2022). A T5-based framework that unifies multiple recommendation tasks into a natural-language generation paradigm. P5 maps each task to a specific text prompt, feeds it into T5, and decodes item IDs as generated tokens. By leveraging T5’s pre-trained language understanding, P5 can flexibly handle diverse recommendation subtasks, though it relies on ID tokens for item representation.
  • BIGRec (Bao et al., 2023a). A LLM-based approach for recommendation that converts a user’s interaction history sequence into a textual prompt. The model generates a natural‐language continuation, and a grounding module maps the generated text back to concrete item.
  • CFT (Zhang et al., 2024b). A method that enhances modeling of user behavior sequences through counterfactual fine-tuning. While CFT improves overall sequence understanding, it does not explicitly encode the temporal ordering of interactions.
5.1.4. Implementation Details
For traditional methods, we use one self-attention blocks in SASRec and one GRU layer in GRU4Rec, other settings follow the original papers. For LLM-based models, we select Llama3-8B (Dubey et al., 2024) as the backbone and use the AdamW (Loshchilov and Hutter, 2019) optimizer with a learning rate of $1\times 10^{-4}$. Specifically, we implement Llama-based P5 in OpenP5 paper (Xu et al., 2024; Hua et al., 2023). For CETRec, we adopt the two position embedding schemes introduced earlier—SinPE and RoPE—as our implementations of temporal embeddings. Besides, we tune the $\lambda$, the hyperparameter controlling the weight of counterfactual loss in our CT task, in the range $\{0.01,0.1,0.5,1,10\}$. In inference stage, we follow the process in CFT paper (Zhang et al., 2024b). We generate five items, finding the most closely matched actual item for each of them to generate the Top-5 recommendation list. The Top-10 recommendation list is generated from the second-best matched items of each item. All experiments are conducted with Python 3.12, PyTorch 2.2.2 and transformers 4.43.4.
Table 4. Ablation results with different positional embedding methods and training variants.
MovieLens Steam
HR@5 NDCG@5 HR@10 NDCG@10 HR@5 NDCG@5 HR@10 NDCG@10
SinPE CETRec 0.1101 0.0693 0.1209 0.0728 0.0670 0.0436 0.0779 0.0471
w.o. CT 0.1035 0.0634 0.1167 0.0676 0.0620 0.0407 0.0720 0.0440
RoPE CETRec 0.1283 0.0779 0.1365 0.0806 0.0888 0.0635 0.0997 0.0670
w.o. CT 0.1192 0.0745 0.1300 0.0781 0.0737 0.0474 0.0821 0.0501
w.o. Both 0.0985 0.0614 0.1093 0.0650 0.0595 0.0408 0.0687 0.0438
5.2. Performance Comparison(RQ1 & RQ2)
We first compare the recommendation performance of baseline methods and CETRec implemented with SinPE and RoPE. The evaluation results are summerized in Table 3. We draw the following observations from the results:
  • CETRec with RoPE outperforms all the baseline models on two datasets except HR@10 on Steam, achiving an average relative improvement of 11.6% across all metrics and datasets. This shows the effectiveness of CETRec.
  • There remains a performance gap between traditional and LLM-based methods. This is partly because traditional models rely on ID embeddings to capture sequential patterns, thereby neglecting the rich semantic information that LLMs leverage for deeper item understanding. Additionally, the orders-of-magnitude difference in parameter scale between traditional models and LLMs contributes to their differing inference capabilities.
  • P5 underperforms most LLM‑based methods across several metrics, primarily because it represents items by their IDs and thus sacrifices the deeper semantic understanding that LLMs can provide. However, as $K$ increases, P5’s gains outpace those of other LLM‑based models. This is explained by its inference strategy: P5 generates a list of 10 items for a Top‑10 recommendation, rather than 5 items in compared methods, which yields greater recommendation diversity. This behavior is consistent with observations reported in the CFT paper (Zhang et al., 2024b).
  • CETRec implemented with RoPE performances better than with SinPE. This discrepancy can be attributed to two factors. First, by Equation 5, RoPE’s effect depends only on the difference between tokens’ position embedding indices $n$ and $m$. In interaction sequences, items that occur earlier receive temporal embedding indices that are farther from the current position, resulting in larger rotation angles and greater deviation from their original representations. This semantic‑order–based encoding of temporal distance is not captured by SinPE. Second, the Llama architecture (Dubey et al., 2024) itself employs RoPE as its positional embedding scheme, enabling a more seamless integration with LLM internals compared to SinPE.
5.3. Ablation Study(RQ3)
To enhance the temporal sensitivity of LLMs for sequential recommendation, CETRec firstly proposes temporal embedding and accordingly designs the counterfactual tuning task. After demonstrating the model’s efficacy, we conduct ablation experiments to validate the effectiveness of its two core components. We introduce two variants:
  • w.o. CT disables the counterfactual tuning and only allows tuning by the temporal-aware loss $L_{TA}$ in Equation 9, this is equivalent to setting $\lambda$ to 0 in Equation 10.
  • w.o. Both variant disables both components. Since counterfactual tuning depends on the presence of temporal embeddings, we cannot disable temporal embeddings independently. However, we can recognize the effect of temporal embedding via camparing the two variants. Notably, under our experimental setup, this variant is effectively equivalent to the vanilla BIGRec method.
The results are detailed in Table 4, from which we can observe:
  • Disabling counterfactual tuning (w.o. CT) results in a significant performance drop across all datasets and metrics, confirming that this tuning method strengthens the LLM’s understanding of user interaction sequences.
  • The performance changes of w.o. Both variant explains the gap between BIGRec and CETRec. Compared to the results of w.o. CT, the differences prove the effectiveness of temporal embedding.
  • From these performance differences, we conclude that—whether using SinPE or RoPE—both the temporal embedding module and the counterfactual tuning component are critical for CETRec.
Figure 4. The performance change rate of each methods testing on reversed sequences in test set of MovieLens datasets.
[…]
Results to illustrate models’ temporal sensitivity.
5.4. Temporal Sensitivity Analysis(RQ4)
We further conduct experiments to investigate the temporal sensitivity of CETRec. By reversing the interaction sequences, we alter both the absolute order and the relative order of each sequence, thereby changing the temporal information they contain. Therefore, we evaluate the model’s performance on the reversed interaction sequences in the test set and quantify its temporal sensitivity by measuring the change rate to its performance on the original test set. We use the MovieLens dataset as an example. The results are shown in Figure 4, and we draw the following conclusions:
  • Both CETRec and the w.o. CT variant increase the performance change rate, demonstrating the effectiveness of both counterfactual tuning and temporal embeddings. However, CETRec’s additional gain over the w.o. CT variant is smaller. This is partly because counterfactual tuning builds on the sequence order captured by temporal embeddings, and partly because, in a recommendation sequence composed of both items and temporal signals, the latter contribute only so much—so the change rate cannot grow without bound.
  • CFT’s change rate is larger than BIGRec’s but still lower than CETRec’s. This is because CFT improves the model’s understanding of the overall interaction sequence, but its temporal modeling granularity is not as fine‐grained as that of CETRec.
  • CETRec with SinPE exhibits a higher change rate than with RoPE. This is because SinPE, as an absolute position encoding, induces larger and more consistent shifts in token embeddings, causing the model to react more strongly to differences in temporal embeddings and thus demonstrate greater temporal sensitivity. This also shows that higher temporal sensitivity does not necessarily translate into better recommendation performance.
6. Conclusion
In this paper, we identify and analyze a critical limitation of LLM-based sequential recommendation systems: their inability to fully leverage the rich temporal information embedded in user interaction sequences. This limitation stems from fundamental architectural constraints in Transformer-based models, where self-attention mechanisms lack inherent sequence ordering and position embeddings designed for natural language prove inadequate for modeling user preference evolution. To address this challenge, we propose Counterfactual Enhanced Temporal Framework for LLM‑Based Recommendation (CETRec), a novel approach grounded in causal inference principles. CETRec first conceptualizes temporal order as an independent causal factor in recommendation outcomes, distinct from item content. By applying temporal embeddings at the item level and constructing a counterfactual world where temporal information is erased, we design a specialized tuning objective that directly optimizes the model’s temporal sensitivity. Our experiments across two real-world datasets demonstrate that CETRec consistently outperforms state-of-the-art baselines. Ablation studies confirm that both temporal embedding and counterfactual tuning components contribute significantly to performance gains, while our temporal sensitivity analysis validates CETRec’s enhanced ability to capture both absolute and relative ordering information.
Looking forward, this work opens several promising research directions. While our current implementation assigns temporal embeddings based solely on interaction order, future work could explore more sophisticated strategies—such as incorporating actual time intervals between interactions, modeling seasonal patterns in user behavior, or integrating collaborative filtering signals—to better capture the complex temporal dynamics in recommendation scenarios. Additionally, extending our counterfactual inference framework to other aspects of recommendation systems, such as user context or item popularity, could yield further improvements in recommendation quality and explainability.
7. GenAI Usage Disclosure
LLM is used to polish the writing of the paper.